OpenStudy (blondey_315):

Solve for n:|3n+5|=7

myininaya (myininaya):

|f(x)|=7 implies f(x)=7 or f(x)=-7

myininaya (myininaya):

You have two equations to solve

OpenStudy (anonymous):

|3n+5|=7 (3n+5)^2 = 7^2 3n + 5 = sqrt(7^2) 3n+5 = 7 3n = 7-5 3n = 2 n = 2/3

OpenStudy (anonymous):

It's like this 3n + 5 = 7 -5 -5 3n = 2 Divide both by 3 so "n" is alone n=2/3

myininaya (myininaya):

You have two equations to solve....

myininaya (myininaya):

3n+5=7 or 3n+5=-7

myininaya (myininaya):

You could do it the way @gerryliyana started off ... square both sides... but you will get two solutions not one

OpenStudy (anonymous):

or n = -4

OpenStudy (anonymous):

i've not done anyway

myininaya (myininaya):

\[(3n+5)^2=7^2 \\ (3n+5)(3n+5)=49 \\ 9n^2+15n+15n+25=49 \\ 9n^2+30n+25=49 \\ 9n^2+30n+25-49=0 \\ 9n^2+30n-24=0 \\ 3n^2+10n-8 =0\text{ divided both sides by 3 } \\ 3(-8)=(-12)(2)=(12)(-2) \text{ and } 12+(-2)=10 \\ 3n^2+12n-2n-8=0 \\ 3n(n+4)-2(n+4)=0 \\ (3n-2)(n+4)=0\]

myininaya (myininaya):

then you know where to go from here